OpenStudy (anonymous):

f(x) = 3/1+2e^-5x Find f`(x) and f"(x)

OpenStudy (lalaly):

do you know the quotient or product rule?

OpenStudy (anonymous):

yes but I don`t seem to get the right answer!

OpenStudy (lalaly):

which one you prefer to use?

OpenStudy (anonymous):

The quotient rule

OpenStudy (lalaly):

ok let h(x)=3 and g(x) = 1+2e^-5x \[\large{f'(x)=\frac{h'(x)g(x)-h(x)g'(x)}{g(x)^2}}\]

OpenStudy (lalaly):

so \[\large{f'(x)= \frac{ (0 \times -1+2e^{-5x}) - 3 \times(-10e^{-5x}}{(1+2e^{-5x})^2}}\]

OpenStudy (lalaly):

\[=\large{\frac{30e^{-5x}}{(1+2e^{-5x})^2}}\]

OpenStudy (lalaly):

for the second derivative, u do the same....try doing it and if u get stuck ill help u

OpenStudy (anonymous):

lalaly how did you get a -1 and a -3?

OpenStudy (anonymous):

shouldn't it have been 0 x 1+2e^-5x in the numerator?

OpenStudy (lalaly):

ignore the -1 that got there by mistake

OpenStudy (anonymous):

ok thanks

OpenStudy (lalaly):

ignore the negative sign only not the 1

OpenStudy (anonymous):

yeah I get it